# Contemporary tensor network approaches to gapless and topological phases in an extended Bose-Hubbard ladder

### v1

@@ Abstract
Introduction. The study of novel exotic equilibrium phases is one of the frontiers in many-body physics. The most challenging task is to evade the exponential growth of Hilbert space, while precisely and sufficiently capturing the ground/low-excited state properties. To this end, various numerical methods have been proposed, such as exact diagonalization, quantum Monte Carlo (QMC) Fehske et al. (2008), tensor network algorithms Schollwöck (2011) and machine learning inspired approaches  Dawid et al. (2023). In particular, the tensor network (TN) approach provides access to large system sizes without model-dependent limitations, while capturing relevant correlations at a reasonable computational cost. The celebrated density matrix renormalization group (DMRG) White (1992) is a leading example of this class of methods wherein low energy states are obtained effectively as variationally optimised matrix product states (MPS). Extended TN methods specialized for specific motivations have been proposed, such as time-evolving block decimation (TEBD)Vidal (2003, 2004) for (real/imaginary) time evolution, projected entangled pair states (PEPS)Verstraete and Cirac (2004); Orús (2014) for 2D systems, as well as the multiscale entanglement renormalization ansatz (MERA) Vidal (2007, 2008) for critical systems. The TN method also provides access to infinite systems where the size/boundary effects can be discarded. The state-of-the-art infinite TN is the variational uniform matrix product state (VUMPS) Zauner-Stauber et al. (2018), which eliminates the disadvantages of conventional methods, such as the infinite DMRG/TEBD (iDMRG/iTEBD) Vidal (2003, 2007); McCulloch (2008); namely it directly deals with infinite systems by variationally optimizing uniform MPS.
@@
Interesting target phases for TNs are topological phases Senthil (2015); Wen (2017) in interacting 1D/quasi-1D/2D systems. The simplest example of a topological phase is the so-called symmetry-protected-topological (SPT) phase Senthil (2015) Despite significant advances in the study of non-interacting SPT Schnyder et al. (2008); Kitaev (2009); Bernevig and Hughes (2013), a comprehensive understanding of interacting/interaction-induced SPT Rachel (2018) is still challenging, remaining a hot topic in modern physics. TN approaches facilitate access to crucial quantities characterizing interacting SPT phases, such as the non-local string order Oshikawa (1992), degenerate edge modes and the degeneracy in the entanglement spectrum Li and Haldane (2008); Pollmann et al. (2010, 2012), which can be difficult to access with other methods, such as the QMC.
@@
Interacting SPT phases are hosted by the 1D Bose-Hubbard (BH) model with non-local interactions Góral et al. (2002); Sengupta et al. (2005); Batrouni et al. (2006); Dalla Torre et al. (2006); Mishra et al. (2009); Iskin (2011); Kimura (2011); Rossini and Fazio (2012); Ohgoe et al. (2012); Bogner et al. (2019). This model has a richer phase diagram than the standard BH model Jaksch et al. (1998) where the competition between hopping of bosonic particles and on-site repulsion leads to a quantum phase transition only between a superfluid (SF) phase and a Mott-insulating (MI) phase, observed experimentally with cold atoms in optical lattices  Greiner et al. (2002). In the extended BH model Dutta et al. (2015), recently realized with dipolar excitons on a synthetic square lattice  Lagoin et al. (2022) and with dipolar atoms Su et al. (2023), other insulating phases appear, including insulators with density wave (DW) modulations, and in 1D an insulating SPT phase, known as a Haldane insulator (HI) Dalla Torre et al. (2006) characterized by nonlocal correlations. In contrast to the DW insulator, the HI phase respects the discrete translational symmetry of the lattice. Nevertheless, a peculiar DW-like pattern of alternating high- and low-density sites is present, hidden by strings of arbitrary numbers of sites with average density. This hidden string ordered phase is the bosonic analog of the spin-1 Haldane phase and is an interacting SPT phase similarly characterized by localized gapless edge excitations and degeneracy in the entanglement spectrum Ejima et al. (2014).
@@
An interesting question concerns the possible coexistence of features of these insulators, such as DW or string order, and superfluid correlations. In the case of DW order, the coexistence with superfluid correlations gives rise to an exotic phase known as the supersolid phase, which has recently been observed with cold atoms Guo et al. (2019); Norcia et al. (2021); Chisholm et al. (2024). The coexistence of superfluidity and Haldane string order has only been reported theoretically in Ref. Lv and Wang, 2016, where Monte-Carlo simulation suggests that such a Haldane superfluid (HSF) arises by coupling two BH chains with nearest-neighbor interactions in a parameter regime where the independent chains would be in the SF phase. Such two-leg ladders have recently been studied also from the point of view of supersolidity at arbitrary filling factors Watanabe et al. (2024), and novel topological phases Wellnitz et al. (2024); Korbmacher et al. (2025). Experimentally, a cold atom Fermi-Hubbard two-leg ladder has allowed for realizing the string order in the Haldane phase Sompet et al. (2022).
@@
However, from both experimental and theoretical points of view, the precise determination of the phase diagram is extremely challenging, as the boundaries of string order and quasi-long-range SF correlations in 1D systems are washed out through finite-size effects. It raises a question regarding the existence of the HSF phase reported by the QMC simulation Lv and Wang (2016), which is supported only by the scanty string order. Therefore, this Letter revisits the scenario using TN methods for (quasi-)1D finite/infinite systems. Specifically, we use the DMRG and the VUMPS methods for finite/infinite systems using the ITensor library Fishman et al. (2022) and primarily focus on obtaining the precise transition points by studying non-local correlations. We show that only the VUMPS calculations provide very sharp boundaries, which are crucial to accurately determine the different phases of the model. While the DMRG results exhibit regimes with overlapping string and SF order, providing room for the putative Haldane superfluid (pHSF) phase, the string order in the VUMPS calculation sharply drops to zero in the parameter regime where superfluid correlations appear in the system. This finding showcases the crucial advantage of this infinite-system tensor network method in characterizing topological and critical phases. The investigations of the topological properties regarding the edge mode and the entanglement spectrum (provided in the supplemental material) further support the VUMPS results.
@@
where, $\hat{\alpha}_{i}(\hat{\alpha}_{i}^{\dagger})$ is annihilation (creation) operator acting on $i$-th site either ladder $a$ or $b$, and the particle number operator is defined as $\hat{n}^{\alpha}_{i}=\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i}(\alpha=a,b)$. We consider the unit-filling case where $N^{\alpha}\equiv\sum_{i}\Braket{\hat{n}_{i}^{\alpha}}=L\ (\alpha=a,b)$, with $L$ being the length of the ladder. This is a conserved quantity, as we assume no hopping between the legs, but only between neighboring sites along each leg, with hopping amplitude $J$. In each leg, we consider the on-site interaction with strength $U$ and the nearest-neighbor interaction with strength $V$. The particles on the same rung but in different legs interact via the inter-leg interaction $U_{ab}$. Here, we are particularly interested in the Haldane superfluid phase, which we are targeting using the same system parameters as previous work Lv and Wang (2016). Specifically, in both legs we choose $U=1.0$ and $V=0.75$, and tune $J$ in order to reach different phases. A small interaction parameter $U_{ab}=0.04$ connects the two legs.
@@
The HI arises when local and non-local interactions are comparable. It is a bulk gapped phase where the DW order described above is diluted (or "hidden") by strings of consecutive sites populated with average density. This phase resembles the topologically ordered state of spin-1 chains. Other characteristic features of the HI phase are degenerate edge states, and non-trivial degeneracy in the entanglement spectrum Li and Haldane (2008); Pollmann et al. (2010, 2012). The most straightforward and well-used quantity for the characterization of the HI is the so-called string order given by
@@ Results.
We first calculate the ground state properties in finite systems using the DMRG method. In the DMRG calculation, we keep up to $800$ states and require the energy and entropy precisions as $1.0\times 10^{-8}$ and $1.0\times 10^{-5}$, respectively. In this paper, we consider open boundary conditions and apply a bias potential, breaking the (edge state’s) degeneracy in the DW (HI) phase. To this end, we apply additional on-site energy $\epsilon^{\alpha}_{i},\alpha\in\{a,b\}$ at the edges $i=1,L$. As we consider the repulsive on-site interaction $U_{ab}$ between two legs, it is natural to choose the bias potential to pin the particles alternatively in each leg, namely, $\epsilon^{a}_{1}=\epsilon^{b}_{L}=0.1,\epsilon^{a}_{L}=\epsilon^{b}_{1}=2.0$. The top panel (a) in Fig. 2 shows how the ground state properties change depending on the hopping for system sizes $L=64,128,256$. The orange circle, blue rectangle, and cyan triangle lines depict the DW, the string order, and finite size SF order parameters, respectively. The system undergoes the transition from the insulating phase to the SF phase depending on the hopping. When the hopping is small, the DW and string order parameters are finite, and the system is in the DW phase. As the hopping increases, the DW order parameter drops to zero while the string order parameter is finite, showing the transition to the HI phase. Even in the SF regime, the string order is still finite, and the system holds the superfluidity and the finite string order parameter simultaneously. This phase fulfills the definition of the HSF phase, claimed in Ref. Lv and Wang, 2016. We show the SF correlation function at $L=256$ along the HI, the pHSF phases ($J=0.3\sim 0.36$) in panel (b) in Fig. 2. The tail of the SF correlation function shows the sharper fall in the HI (gray) compared to he pHSF (red) phase, indicating the pHSF regimes exhibit a long-range correlation. Panel (c) in Fig. 2 represents the system size scaling of the string order parameter in the pHSF ($J=0.38$) and the SF ($J=0.46$) phases. The string order parameter is close to zero at the infinite system size limit in the SF, while it converges to a finite value in the pHSF regime. These results are consistent with the QMC calculations (at small finite temperature) with periodic boundary condition Lv and Wang (2016).
@@
Numerical methods for finite systems, such as the DMRG and the QMC, require system size scaling and the artificial choices of boundary conditions for computational accuracy and stability. To eliminate these needs, one can consider using periodic boundary conditions (PBC); however, the PBC should usually be avoided in terms of the DMRG due to its high computational costs. Another choice is to deal with infinite systems using specialized TN methods, such as the infinite DMRG (iDMRG) and the VUMPS (see supplemental materials for more details). The VUMPS globally optimizes and updates the entire system and directly addresses the thermodynamic limit at each iteration without breaking translation invariance, achieving convergence with fewer iterations than the iDMRG, especially for phases with long correlations. The superior performance of the VUMPS to capture the long-range correlation over the iDMRG is also shown in the conventional BH model Kiely and Mueller (2022). Notably, the VUMPS works efficiently in systems with long-range interactions Zauner-Stauber et al. (2018), which is essentially important in quasi-1D systems, where effective long-range interactions are introduced by mapping quasi-1D structures into 1D chains. For these advantages, we adopt the VUMPS in the following to obtain a more precise phase boundary, especially between the HI and the pHSF, by calculating the SF and the string correlations in infinite systems.
@@
In Fig. 3, we plot the SF and the string order parameters across the HI and the pHSF phases with the maximum bond dimension $\chi=250,450$ in panels (a) and (b), respectively. The string order parameter in the infinite system is calculated as follows: (1) we first calculate the string order parameter in the finite systems, which are composed of copies of the unit cell obtained via the VUMPS algorithm, (2) Then, we conduct the system size scaling to obtain the string order in an infinite system. As examples, we show the string correlation function and the system size scaling of the string order parameter in a window of hopping values suggested, by our finite DMRG shown in Fig.2 and previous QMC results Lv and Wang (2016), to host the phase boundary of the HI and the pHSF in panels (d) and (e) in Fig. 3. As in the case of the DMRG calculation, we define the superfluidity indicator as values of the SF correlation function Eq. (2) at a certain two-point distance. Here, we use $R=200$, namely,
@@
Conclusion. We have investigated the phase transition of an extended BH ladder near the regime where a pHSF of coexisting SF and string orders has been reported Lv and Wang (2016). Using the DMRG and the VUMPS method, we have calculated the string and the SF correlations in finite and infinite systems. While the DMRG calculation, even after finite-size scaling, shows an overlap regime of two correlations, the string order sharply vanishes in the VUMPS calculation at the boundary where the SF correlations significantly appear, leaving no room for the pHSF phase. With further investigation of the topological aspects provided in the supplemental material, we conclude that pHSF is, in fact, an ordinary SF phase, contrary to the finite-size DMRG results and Monte Carlo results from Ref. Lv and Wang, 2016. Although our numerical study does not rule out the existence of the HSF phase in other parameter regimes, we speculate that the gapless nature of a superfluid is not compatible with the SPT order of the Haldane phase. Our investigation showcases the advantages of infinite-size tensor network methods for studying topological and gapless phases, as compared to finite-sized scaled DMRG or Monte Carlo methods.
@@ I.1.1 Algorithm
Here, we briefly summarize the algorithm of the two TN methods used in the main text, the DMRG and the VUMPS. We refer the reader Refs. White (1992); Schollwöck (2011) for the DMRG and  Zauner-Stauber et al. (2018); Vanderstraeten et al. (2019) for the VUMPS for the further detail.
@@
The bond dimension $\chi$ does not increase during the VUMPS algorithm, as it is based on a one-site DMRG-like scheme. However, a subspace expansion technique can be used to obtain the desired bond dimension, which is well-explained in Zauner-Stauber et al. (2018). The algorithm repeats these optimization processes until it meets the criteria regarding the error given by $\epsilon=\max(\epsilon_{L},\epsilon_{R})$.

### v2

@@ Abstract
Introduction. The study of novel exotic equilibrium phases is one of the frontiers in many-body physics. The most challenging task is to evade the exponential growth of Hilbert space, while precisely and sufficiently capturing the ground/low-excited state properties. To this end, various numerical methods have been proposed, such as exact diagonalization, quantum Monte Carlo (QMC) [1], tensor network algorithms [2] and machine learning inspired approaches  [3]. In particular, the tensor network (TN) approach provides access to large system sizes without model-dependent limitations, while capturing relevant correlations at a reasonable computational cost. The celebrated density matrix renormalization group (DMRG) [4] is a leading example of this class of methods wherein low energy states are obtained effectively as variationally optimised matrix product states (MPS). Extended TN methods specialized for specific motivations have been proposed, such as time-evolving block decimation (TEBD)[5, 6] for (real/imaginary) time evolution, projected entangled pair states (PEPS)[7, 8] for 2D systems, as well as the multiscale entanglement renormalization ansatz (MERA) [9, 10] for critical systems. The TN method also provides access to infinite systems where the size/boundary effects can be discarded. The state-of-the-art infinite TN is the variational uniform matrix product state (VUMPS) [11], which eliminates the disadvantages of conventional methods, such as the infinite DMRG/TEBD (iDMRG/iTEBD) [5, 9, 12]; namely it directly deals with infinite systems by variationally optimizing uniform MPS.
@@
Interesting target phases for TNs are topological phases [13, 14] in interacting 1D/quasi-1D/2D systems. The simplest example of a topological phase is the so-called symmetry-protected-topological (SPT) phase [13] Despite significant advances in the study of non-interacting SPT [15, 16, 17], a comprehensive understanding of interacting/interaction-induced SPT [18] is still challenging, remaining a hot topic in modern physics. TN approaches facilitate access to crucial quantities characterizing interacting SPT phases, such as the non-local string order [19], degenerate edge modes and the degeneracy in the entanglement spectrum [20, 21, 22], which can be difficult to access with other methods, such as the QMC.
@@
Interacting SPT phases are hosted by the 1D Bose-Hubbard (BH) model with non-local interactions [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. This model has a richer phase diagram than the standard BH model [33] where the competition between hopping of bosonic particles and on-site repulsion leads to a quantum phase transition only between a superfluid (SF) phase and a Mott-insulating (MI) phase, observed experimentally with cold atoms in optical lattices  [34]. In the extended BH model [35], recently realized with dipolar excitons on a synthetic square lattice  [36] and with dipolar atoms [37], other insulating phases appear, including insulators with density wave (DW) modulations, and in 1D an insulating SPT phase, known as a Haldane insulator (HI) [26] characterized by nonlocal correlations. In contrast to the DW insulator, the HI phase respects the discrete translational symmetry of the lattice. Nevertheless, a peculiar DW-like pattern of alternating high- and low-density sites is present, hidden by strings of arbitrary numbers of sites with average density. This hidden string ordered phase is the bosonic analog of the spin-1 Haldane phase and is an interacting SPT phase similarly characterized by localized gapless edge excitations and degeneracy in the entanglement spectrum [38].
@@
An interesting question concerns the possible coexistence of features of these insulators, such as DW or string order, and superfluid correlations. In the case of DW order, the coexistence with superfluid correlations gives rise to an exotic phase known as the supersolid phase, which has recently been observed with cold atoms [39, 40, 41]. The coexistence of superfluidity and Haldane string order has only been reported theoretically in Ref. 42, where Monte-Carlo simulation suggests that such a Haldane superfluid (HSF) arises by coupling two BH chains with nearest-neighbor interactions in a parameter regime where the independent chains would be in the SF phase. Such two-leg ladders have recently been studied also from the point of view of supersolidity at arbitrary filling factors [43], and novel topological phases [44, 45]. Experimentally, a cold atom Fermi-Hubbard two-leg ladder has allowed for realizing the string order in the Haldane phase [46].
@@
However, from both experimental and theoretical points of view, the precise determination of the phase diagram is extremely challenging, as the boundaries of string order and quasi-long-range SF correlations in 1D systems are washed out through finite-size effects. It raises a question regarding the existence of the HSF phase reported by the QMC simulation [42], which is supported only by the scanty string order. Therefore, this Letter revisits the scenario using TN methods for (quasi-)1D finite/infinite systems. Specifically, we use the DMRG and the VUMPS methods for finite/infinite systems using the ITensor library [47] and primarily focus on obtaining the precise transition points by studying non-local correlations. We show that only the VUMPS calculations provide very sharp boundaries, which are crucial to accurately determine the different phases of the model. While the DMRG results exhibit regimes with overlapping string and SF order, providing room for the putative Haldane superfluid (pHSF) phase, the string order in the VUMPS calculation sharply drops to zero in the parameter regime where superfluid correlations appear in the system. This finding showcases the crucial advantage of this infinite-system tensor network method in characterizing topological and critical phases. The investigations of the topological properties regarding the edge mode and the entanglement spectrum (provided in the supplemental material) further support the VUMPS results.
@@
where, $\hat{\alpha}_{i}(\hat{\alpha}_{i}^{\dagger})$ is annihilation (creation) operator acting on $i$-th site either ladder $a$ or $b$, and the particle number operator is defined as $\hat{n}^{\alpha}_{i}=\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i}(\alpha=a,b)$. We consider the unit-filling case where $N^{\alpha}\equiv\sum_{i}\Braket{\hat{n}_{i}^{\alpha}}=L\ (\alpha=a,b)$, with $L$ being the length of the ladder. This is a conserved quantity, as we assume no hopping between the legs, but only between neighboring sites along each leg, with hopping amplitude $J$. In each leg, we consider the on-site interaction with strength $U$ and the nearest-neighbor interaction with strength $V$. The particles on the same rung but in different legs interact via the inter-leg interaction $U_{ab}$. Here, we are particularly interested in the Haldane superfluid phase, which we are targeting using the same system parameters as previous work [42]. Specifically, in both legs we choose $U=1.0$ and $V=0.75$, and tune $J$ in order to reach different phases. A small interaction parameter $U_{ab}=0.04$ connects the two legs.
@@
The HI arises when local and non-local interactions are comparable. It is a bulk gapped phase where the DW order described above is diluted (or "hidden") by strings of consecutive sites populated with average density. This phase resembles the topologically ordered state of spin-1 chains. Other characteristic features of the HI phase are degenerate edge states, and non-trivial degeneracy in the entanglement spectrum [20, 21, 22]. The most straightforward and well-used quantity for the characterization of the HI is the so-called string order given by
@@ Results.
We first calculate the ground state properties in finite systems using the DMRG method. In the DMRG calculation, we keep up to $800$ states and require the energy and entropy precisions as $1.0\times 10^{-8}$ and $1.0\times 10^{-5}$, respectively. In this paper, we consider open boundary conditions and apply a bias potential, breaking the (edge state’s) degeneracy in the DW (HI) phase. To this end, we apply additional on-site energy $\epsilon^{\alpha}_{i},\alpha\in\{a,b\}$ at the edges $i=1,L$. As we consider the repulsive on-site interaction $U_{ab}$ between two legs, it is natural to choose the bias potential to pin the particles alternatively in each leg, namely, $\epsilon^{a}_{1}=\epsilon^{b}_{L}=0.1,\epsilon^{a}_{L}=\epsilon^{b}_{1}=2.0$. The top panel (a) in Fig. 2 shows how the ground state properties change depending on the hopping for system sizes $L=64,128,256$. The orange circle, blue rectangle, and cyan triangle lines depict the DW, the string order, and finite size SF order parameters, respectively. The system undergoes the transition from the insulating phase to the SF phase depending on the hopping. When the hopping is small, the DW and string order parameters are finite, and the system is in the DW phase. As the hopping increases, the DW order parameter drops to zero while the string order parameter is finite, showing the transition to the HI phase. Even in the SF regime, the string order is still finite, and the system holds the superfluidity and the finite string order parameter simultaneously. This phase fulfills the definition of the HSF phase, claimed in Ref. 42. We show the SF correlation function at $L=256$ along the HI, the pHSF phases ($J=0.3\sim 0.36$) in panel (b) in Fig. 2. The tail of the SF correlation function shows the sharper fall in the HI (gray) compared to he pHSF (red) phase, indicating the pHSF regimes exhibit a long-range correlation. Panel (c) in Fig. 2 represents the system size scaling of the string order parameter in the pHSF ($J=0.38$) and the SF ($J=0.46$) phases. The string order parameter is close to zero at the infinite system size limit in the SF, while it converges to a finite value in the pHSF regime. These results are consistent with the QMC calculations (at small finite temperature) with periodic boundary condition [42].
@@
Numerical methods for finite systems, such as the DMRG and the QMC, require system size scaling and the artificial choices of boundary conditions for computational accuracy and stability. To eliminate these needs, one can consider using periodic boundary conditions (PBC); however, the PBC should usually be avoided in terms of the DMRG due to its high computational costs. Another choice is to deal with infinite systems using specialized TN methods, such as the infinite DMRG (iDMRG) and the VUMPS (see supplemental materials for more details). The VUMPS globally optimizes and updates the entire system and directly addresses the thermodynamic limit at each iteration without breaking translation invariance, achieving convergence with fewer iterations than the iDMRG, especially for phases with long correlations. The superior performance of the VUMPS to capture the long-range correlation over the iDMRG is also shown in the conventional BH model [48]. Notably, the VUMPS works efficiently in systems with long-range interactions [11], which is essentially important in quasi-1D systems, where effective long-range interactions are introduced by mapping quasi-1D structures into 1D chains. For these advantages, we adopt the VUMPS in the following to obtain a more precise phase boundary, especially between the HI and the pHSF, by calculating the SF and the string correlations in infinite systems.
@@
In Fig. 3, we plot the SF and the string order parameters across the HI and the pHSF phases with the maximum bond dimension $\chi=250,450$ in panels (a) and (b), respectively. The string order parameter in the infinite system is calculated as follows: (1) we first calculate the string order parameter in the finite systems, which are composed of copies of the unit cell obtained via the VUMPS algorithm, (2) Then, we conduct the system size scaling to obtain the string order in an infinite system. As examples, we show the string correlation function and the system size scaling of the string order parameter in a window of hopping values suggested, by our finite DMRG shown in Fig.2 and previous QMC results [42], to host the phase boundary of the HI and the pHSF in panels (d) and (e) in Fig. 3. As in the case of the DMRG calculation, we define the superfluidity indicator as values of the SF correlation function Eq. (2) at a certain two-point distance. Here, we use $R=200$, namely,
@@
Conclusion. We have investigated the phase transition of an extended BH ladder near the regime where a pHSF of coexisting SF and string orders has been reported [42]. Using the DMRG and the VUMPS method, we have calculated the string and the SF correlations in finite and infinite systems. While the DMRG calculation, even after finite-size scaling, shows an overlap regime of two correlations, the string order sharply vanishes in the VUMPS calculation at the boundary where the SF correlations significantly appear, leaving no room for the pHSF phase. With further investigation of the topological aspects provided in the supplemental material, we conclude that pHSF is, in fact, an ordinary SF phase, contrary to the finite-size DMRG results and Monte Carlo results from Ref. 42. Although our numerical study does not rule out the existence of the HSF phase in other parameter regimes, we speculate that the gapless nature of a superfluid is not compatible with the SPT order of the Haldane phase. Our investigation showcases the advantages of infinite-size tensor network methods for studying topological and gapless phases, as compared to finite-sized scaled DMRG or Monte Carlo methods.
@@ I.1.1 Algorithm
Here, we briefly summarize the algorithm of the two TN methods used in the main text, the DMRG and the VUMPS. We refer the reader Refs. [4, 2] for the DMRG and  [11, 49] for the VUMPS for the further detail.
@@
The bond dimension $\chi$ does not increase during the VUMPS algorithm, as it is based on a one-site DMRG-like scheme. However, a subspace expansion technique can be used to obtain the desired bond dimension, which is well-explained in [11]. The algorithm repeats these optimization processes until it meets the criteria regarding the error given by $\epsilon=\max(\epsilon_{L},\epsilon_{R})$.